OpenStudy (anonymous):

find the y-intercept of 6x-y=(-12)

OpenStudy (ghazi):

divide whole equation by -12 and see what comes under y

OpenStudy (anonymous):

to find y intercept put x= 0 then you will y = 12

OpenStudy (anonymous):

so positive 1/12 ?

OpenStudy (ghazi):

no..intercept is 12

OpenStudy (ghazi):

it's X/a + Y/b= 1 where a and b are intercepts

OpenStudy (ghazi):

is it clear?

OpenStudy (anonymous):

a little used form for the equation of a line is called the intercept form (this is different from slope-intercept form). it is written in this form: \(\large \frac{x}{a}+\frac{y}{b}=1 \) where a and b are the x and y intercepts, respectively.